OpenStudy (anonymous):

The city of Valley Grove is considering shortening the length of the school day. The school board hired Mr. Kent to do a survey to help them decide what to do. a. Mr. Kent interviewed 54 students as they left the school. Is this a random sample? Explain. b. Mr. Kent used this survey question: “Should the lengthy school day, which now extends for 7.5 hours, be shortened to 6 hours?” Is this question biased or fair? Explain. c. Of the 54 responses he received, 51 were “yes.” The rest were “no.” What percent of the responses were “no”?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

a) Not necessarily. If he interview the first 54 to leave, he is getting the students who don't participate in after-class activities. Converse for last 54 students. If he selected 54 at random times over a period of several hours, the selection may be random. b) No. Leading question. c) 54 - 51 = 3 3/54 = 0.0555... 0.0555...(100) = 5.6% to 2 significant figures ^ My answer | |

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

is my answer right?

OpenStudy (anonymous):

i tgink it involves testing of hypothesis ..

OpenStudy (anonymous):

@urkhada was my answer right?

OpenStudy (urkhada):

I am on mobile, i don' know.. If I am on computer I can tell u by doing it

OpenStudy (anonymous):

so... my answer is wrong?

OpenStudy (anonymous):

here I updated me answer a. no, becuase Mr. Kent just interviewed students, not others in the Valley Grove. b. biased, becuase he calles it a lengthy school day and lengthy implies that it should be shortened. c. 54-51=3 students who said no 3 (students who said no)/ 54 (total students) 0.055555555555555555555555+ *100 (to turn the decimal into a percent) about 05.5 %

OpenStudy (anonymous):

@sourwing is my answer correct? a. no, becuase Mr. Kent just interviewed students, not others in the Valley Grove. b. biased, becuase he calles it a lengthy school day and lengthy implies that it should be shortened. c. 54-51=3 students who said no 3 (students who said no)/ 54 (total students) 0.055555555555555555555555+ *100 (to turn the decimal into a percent) about 05.5 %